Suppose Q=k^.6L^.6 and that labor and capital owners are paid in units of output. Let input owners be paid the marginal product of the last unit of input purchased.

w=MPL

R=MPK

Will the firm produce enough units to pay it's factors?
Answer #1

Q= k^0.6*L^0.6, Now MPK = 0.6k^-0.4*L^0.6 = 0.6L^0.6/K^0.4, similarly MPL = 0.6K^0.6/L^0.4, If K =4 and L= 4 then we can calculate Q= 6^0.6*6^0.6= 8.58, If K=4, and L =4 then MPK = 0.6*4^0.6/4^0.4 = 0.7917 , Now MPL = 0.6*4^0.6/4^0.4 = 0.7917, Therefore MPL + MPK = 0.7917+0.7917 = 1.5834. It is less than total output. So the firm produce enough units (8.58) to pay its factors.